4. (Sec. 7.2) In a survey of 2004 American adults, 501 of them said that they believed in astrology (a) Calculate and interpret a confidence interval at the 95% confidence level for the proportion of all adult American adults who believe in astrology. (b) Calculate and interpret a 95% lower confidence bound for the proportion of all adult American adults who believe in astrology.
Answer #1

Ans:

Given that

sample proportion=501/2004=0.25

n=2004

a)z value(two tailed) for 95% confidence interval is 1.96

95% confidence interval for p

=0.25+/-1.96*sqrt(0.25*(1-0.25)/2004)

=0.25+/-0.019

=(0.231, 0.269)

b)For 95% lower bound, z value will be -1.645

lower bound for p=0.25-1.645*sqrt(0.25*(1-0.25)/2004)

=0.25-0.016

=0.234